OpenStudy (anonymous):

Determine the critical points, extremal points and intervals of monotony for the following functions on \(\mathbb{R}\): a) \(f(x) = x^{4}+2x^{3}-2x^{2}+1\)

OpenStudy (anonymous):

what's an interval of monotony?

OpenStudy (anonymous):

hmm i dont know maybe its false translation, i ask a friend who can speak german and say you..

OpenStudy (anonymous):

hang on... i'll google it...

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

it seems that an interval of monotony is just where the function is increasing/decreasing...

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

i thought it would be a false translation of my mentor but its ok i thin as you said

OpenStudy (kinggeorge):

So then the question is asking for critical points, max/mins, and the intervals between the critical points.

OpenStudy (anonymous):

and how we start ?

OpenStudy (kinggeorge):

To find these, just take the derivative of the function. Can you do that yourself?

OpenStudy (anonymous):

unfortunately not, but the year before last year i was able to

OpenStudy (anonymous):

a friend of me tought me, and missed to pass this same exam with just 2 points :)

OpenStudy (kinggeorge):

Just go back to the power rule I was describing earlier. So the derivative of \(x^n\) is \(nx^{n-1}\). Using this, the derivative of \(x^4\) is \(4x^3\). Using the same pattern what's the derivative of \(2x^3\)?

OpenStudy (anonymous):

ok i remember

OpenStudy (anonymous):

\[6x^{2}\]

OpenStudy (anonymous):

is that?

OpenStudy (kinggeorge):

Perfect. So now, what's the derivative of \[x^4+2x^3-2x^2+1\]

OpenStudy (anonymous):

\[4x^{3}+6x^{2}-4x+1\]

OpenStudy (anonymous):

can it be?

OpenStudy (kinggeorge):

Just about. You should have\[4x^3+6x^2-4x\]Now that you have this, you need to find the critical points. To do this, set the derivative equal to 0, and solve for \(x\). That means we have\[4x^3+6x^2-4x=x(4x^2+6x-4)=0\]

OpenStudy (kinggeorge):

If we continue factoring it, we get to \[2x(x+2)(2x-1)\]So this has zeroes at \[x=0,\quad x=-2,\quad x=\frac{1}{2}\]These are your critical points.

OpenStudy (anonymous):

why we found here \[x(4x^2+6x-4)=0\]

OpenStudy (anonymous):

i mean why we found this equation zero?

OpenStudy (kinggeorge):

By definition, where the derivative is 0, is where the critical points are. We just looked at that because that's where the critical points are defined.

OpenStudy (anonymous):

ok...

OpenStudy (kinggeorge):

To find the extrema, plug the solutions of that back into the original equation. This would get you \[0^{4}+2(0)^{3}-2(0)^{2}+1=1\]\[(-2)^{4}+2(-2)^{3}-2(-2)^{2}+1=-7\]\[(1/2)^{4}+2(1/2)^{3}-2(1/2)^{2}+1=\frac{11}{16}\](I'm not entirely positive I did the math right on the last one)

OpenStudy (kinggeorge):

These are the y-values of the function at the critical points. That means your extrema occur at the points \[(0, 1),\quad(-2, -7),\quad\left(\frac{1}{2}, \frac{11}{16}\right)\]

OpenStudy (anonymous):

ok George, and is this our end solution ?

OpenStudy (kinggeorge):

Now we just need to find the intervals of monotony. These are the intervals between the \(x\) values. So the intervals of monotony are \[(-\infty,-2),\quad(-2,0),\quad\left(0,\frac{1}{2}\right),\quad\left(\frac{1}{2},\infty\right)\]

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

that 11/16 should be 13/16...

OpenStudy (kinggeorge):

I told you I wasn't sure about that :( So wherever I wrote 11/16, replace it with 13/16

OpenStudy (precal):

you should do a sign analysis to determine local max or local min

OpenStudy (kinggeorge):

We're looking for the extremal points, so we need to find the points anyways.

OpenStudy (precal):

when you solve for 0, those values are your critical points

OpenStudy (precal):

have to determine critical points first then determine if they are extrema

OpenStudy (precal):

2nd derivative will help identify points of inflection

OpenStudy (kinggeorge):

Fair enough, but in this case they are all extrema.

OpenStudy (anonymous):

ok thank you guys